OpenStudy (anonymous):

I need help solving this Matrix. [-8+t 0] = [-5 0] [8 -12] [8 -2y-2] I guessed it was t=13, y=5

OpenStudy (anonymous):

\(t=13\) looks good you solved \(-8+t=-5\) right?

OpenStudy (anonymous):

also \(y=5\) is good as well

OpenStudy (anonymous):

Thanks. : )

OpenStudy (gorv):

well we need to equate corresponding value of matrix when we equate them

OpenStudy (anonymous):

yw anyone whose name is Ebullience is cool by me

OpenStudy (gorv):

t not equal 13 check that once

OpenStudy (anonymous):

Eh, then is it 3? e_e That would probably make more sense. . .

OpenStudy (anonymous):

oh good point!

OpenStudy (anonymous):

\[-8+t=-5\\ t=-5+8\]

OpenStudy (anonymous):

Yeah, nice catch Gorv.

OpenStudy (gorv):

mistake could happen loll